MOVIE TICKETS COST $5 FOR ADULTS AND $2 FOR CHILDREN. ON ONE DAY 785 TICKETS WERE SOLD FOR $3280. HOW MANY CHILDREN'S TICKETS WERE SOLD?
Answer by macston(5194)

A=number of adult tickets; C=number of child tickets
A+C=785
A=785-C Use this to substitute for A.
$5A+$2C=$3280 Substitute for A.
$5(785-C)+$2C=$3280
$3925-$5C+$2C=$3280 Subtract $3925 from each side.
-$3C=-$645 Divide each side by -$3
C=215 ANSWER There were 215 child tickets sold.
CHECK:
A=785-C=785-215=570
$5A+$2C=$3280
$5(570)+$2(215)=$3280
$2850+$430=$3280
$3280=$3280